OpenStudy (anonymous):

r/12 = 1 9/10 @chosenmatt

OpenStudy (kamibug):

First turn 1 and 9/10 into an improper fraction. Then multiply both sides by 12 to solve for r. :)

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

hello @KamiBug

OpenStudy (anonymous):

hello @KamiBug

OpenStudy (kamibug):

1 and 9/10 as an improper fraction is 19/10 So now your equation is r/12 = 19/10 Now you need to isolate the r. So to do that we're going to multiply both sides by twelve so that is cancels on the left side of the equal sign. r = 19/10 * 12 So what is 19/10 * 12? :)

OpenStudy (anonymous):

22.8

OpenStudy (kamibug):

Correct. Great job! ^.^

OpenStudy (anonymous):

but thats not an option?

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

Are your options in fraction form?

OpenStudy (anonymous):

yes

OpenStudy (kamibug):

Okay, then we just need to simplify the fraction 288/10. Can you do that? :)

OpenStudy (anonymous):

no i have no ideo how to sry can you show me plz

OpenStudy (kamibug):

We need to divide by the GCF (Greatest Common Factor) of the two numbers. The GCF of 10 and 288 is 2. So we will divide the numerator and denominator by 2. 288/2 = 144 and 10/2 = 5 So your simplified fraction should be 144/5. :) Is that one of the options?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

Oh! Wait, are the options mixed numbers? Lol

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hello @KamiBug

OpenStudy (kamibug):

Haha, there we go. Then all we gotta do is change it to an improper fraction. It should be 28 and 4/5. Please tell me that it's there? xD

OpenStudy (anonymous):

no @KamiBug

OpenStudy (kamibug):

What are the options?

OpenStudy (anonymous):

22 4/5 12 9/10 10 4/5 19/12

OpenStudy (kamibug):

And are you sure you copied the equation from the question correctly?

OpenStudy (anonymous):

yea r/10 = 1 9/10

OpenStudy (kamibug):

That's really weird b/c none of those answers are equal to the value of r according to this equation...Let me tag some people to see if they know...@amistre64 @ganeshie8 @SolomonZelman @Abhisar

OpenStudy (kamibug):

@amistre64

OpenStudy (amistre64):

the question is indeed broken

OpenStudy (amistre64):

maybe try a screeshot of it, or a camera pic?